# Deployability-Centric Infrastructure-as-Code Generation: An LLM-based Iterative Framework

### v1

@@ VI-B RQ2: What are the primary barriers preventing IaCGen from generating deployable IaC templates?
(1) Missing Value. The most prevalent (14.70%) error across models is the deployment error: “An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [xx] must have values.” It happens when LLMs define parameters without providing corresponding values, which is particularly common with the KeyName, VPC, and Subnet parameters. Sometimes, models deliberately leave parameters undefined when information is not explicitly provided in the prompt (e.g., email addresses for SNS notifications). This error appears to be models’ assumption that resources have already existed in the cloud environment, leading them to create placeholders for users to manually fulfill. Claude-3.7 shows the lowest occurrence rate and holds the highest success rate to resolve the issue with the iterative feedback.
This pattern could stem from the prevalence of such template structures in the GitHub repositories as we observe. Also, it typically requires explicit human guidance such as: “You should create the parameters [xx] in the Resources section instead of the Parameters section.” to fix.
@@
(3) Substitute Nothing. “‘Fn::Sub’ is not needed because there are no variables” is the third most common (3.09%) error. This error typically occurs in the UserData attributes of EC2::Instance and EC2::LaunchTemplate, where the models employ “!Sub |||| despite the absence of variables that require substitution. This pattern likely results from large-scale training corpus that contain numerous examples using this function with UserData, leading models to adopt it as a standard practice regardless of necessity.
@@
(4) Unnecessary Whitespace. “Line xx: too many spaces inside brackets” accounts for 2.92% of all errors. This issue disproportionately affects o3-mini, accounting for 75% of all instances across models. The error typically manifests itself in constructs like “!Select [ 0, !GetAZs ‘’ ]” (Note the whitespaces on the left of ‘0’ and ‘]’) instead of the correct ways ofs “!Select [0, !GetAZs ‘’]”. o3-mini consistently inserts extraneous whitespace adjacent to brackets. This error is particularly resistant to correction, often requiring explicit prompt rather than general instructions to resolve.
@@
(5) Arbitrary Default Value. “Parameter validation failed: parameter value xx for parameter name xx does not exist” denotes the error that occurs when models attempt to resolve missing parameter values by providing arbitrary defaults. Claude-3.5, Claude-3.7, and DeepSeek-R1 demonstrate superior resistence in this category, averaging only 6 occurrences per model compared to approximately 27 occurrences for other models. This suggests that these models hold a more comprehensive understanding of the appropriate default values for the AWS resource parameters.
@@
Answer to RQ2: Most errors are identified in the schema checking stage, and the most common error is Missing Value. The top five errors across the three stages account for around 30% of the total failure instances. Different models exhibit a distinct error pattern, and Claude models show the best performance in general.
Figure 7: PassItr@n scores for evaluated models. Transparent lines refer to the result with general and detailed feedback. Solid lines refer to the results additionally with human feedback.
VI-C RQ3: How effective is Human-in-the-loop feedback strategy for IaCGen?
@@
Human-in-the-loop feedback has emerged as a crucial strategy across machine learning and LLM domains [56, 57, 58], demonstrating significant improvements in complex scenarios especially in SE tasks. Enlightened by this, we incorporate human feedback into our framework to explore its capability on tackling unique challenges of IaC template generation. This level of feedback simulates the guidance that a senior cloud engineer might provide to a junior developer during code review, testing whether LLMs can effectively utilize and incorporate direct expert instructions.
@@
To comprehensively evaluate the effectiveness, we include human feedback in addition to the general and detailed feedback mentioned in Section IV-C. Human feedback represents the highest level of assistance in the three-tier feedback hierarchy, providing explicit correction guidance such as: “Add CidrBlock: 10.0.0.0/16 to the VPC resource” or “Replace the AMI ID with placeholder for correct AMI ID value.” A cloud engineer is asked to craft the human feedback based on the ground truth templates, error messages, and online searching. All human feedback is available in our code repository. We assign the same four attempts for the human feedback as the detailed feedback with similar rationale. Therefore, the maximum attempts for each stage (Format Verification, Schema Checking, and Live Deployment) will be ten iterations.
@@
Answer to RQ3: Human feedback provides consistent improvements across all evaluated models, with gains ranging from 2.6% to 37.9% percentage points. All models achieve a performance of over 90% passItr@25.
VI-D RQ4: How trustworthy is the IaC generated by IaCGen?
TABLE IV: User intent performance (%) for various models when tested using the last generated IaC template. Resource stands for resource level intent match and Attribute stands for attribute level intent match.
Model Resource Attribute Resource+Attribute
@@ VI-D RQ4: How trustworthy is the IaC generated by IaCGen?
User Intent Matching. User requirements for IaC templates are often ambiguously specified due to the inherent complexity and technical nature of cloud infrastructure. To evaluate how well the generated IaC templates align with the intended infrastructure requirements, we employ an intent match evaluation. Specifically, based on the DPIaC-Eval benchmark, a cloud engineering expert formulates intent specifications comprising two elements:
  • Required Resources: the core infrastructure components that must be instantiated within the template.
  • Required Attributes: the essential configuration parameters associated with these resources, formalized using Checkov’s YAML-based custom policy format.
Table IV reveal the varying levels of capability between models to match the user intent. We observe a relatively sound capability on generating the required resources, with an average of 58.8% templates that include all specified resources across all models. In contrast, attribute-level matching presents a greater challenge, with models achieving an average of only 40.5% compliance in generating all required attributes with corresponding values. Only 25.2% of the generated templates satisfy both the resource-level and attribute-level requirements across all models. Claude-3.7 demonstrates superior overall performance, yet only 31.4% of its generated IaC templates fully satisfy the user intent.
TABLE V: Security performance (%) of the generated templates at maximum iterations across models.
Model Policy Pass Unfiltered Comp. Filtered Comp.
@@
Security. Security of IaC templates represents a critical concern in cloud computing, as insecure configurations can lead to data breaches, unauthorized access, and compliance violations [12]. To evaluate the security posture of LLM-generated IaC templates that passed deployability evaluation, we perform the security analysis by employing Checkov [59], an popular open source policy-as-code tool for IaC security check. We define three distinct compliance metrics:
  • Policy Pass Rate: The percentage of security policies satisfied across all security policies checked by Checkov.
  • Unfiltered Compliance Rate: The proportion of generated templates that pass all applicable Checkov security policy checks, irrespective of policy applicability.
  • Filtered Compliance Rate: To account for Checkov’s limited policy coverage, this metric excludes templates to which no relevant security policies can be applied.
@@
Table V shows the security compliance of generated IaC templates, and it reveals an alarmingly low security compliance rate overall. In terms of security policies, 75.3% of the examined security policies are passed on average. Surprisingly, at template-level, only 20.2% unfiltered compliance and 8.4% filtered compliance are obtained. This apparent contradiction suggests that, while models generally adhere to most security rules, critical failures in at least one policy consistently invalidate entire templates. DeepSeek-V3 demonstrates the best performance in both unfiltered and filtered compliance (24.2% and 11.5%). Claude-3.5 ranks the lowest (17.6%) in unfiltered compliance and GPT-4o shows the weakest in filtered compliance (6.1%). Further analysis reveals 55.1% of template failures originate from three AWS services: S3 Bucket, Lambda Function, and EC2 SecurityGroup.
These findings indicate a critical threshold effect—while models successfully implement most security requirements ($>$70.0% policy pass), critical failures in 1-2 policies consistently invalidate entire templates. The discrepancy between high policy-level compliance and catastrophic template-level failure highlights fundamental challenges in the ability of LLMs to holistically satisfy security-critical IaC requirements.
@@
Answer to RQ4: The trustworthiness of LLM-generated IaC templates remains moderate. Only 25.2% fully satisfied user intent, with reasonable resource identification (58.8%) but weaker attribute configuration. Security analysis shows that despite a 75.3% policy pass rate, only 8.4% of templates achieved full compliance.
VII Discussion
VII-A Generalizability of IaCGen
We further test the generalizability of IaCGen on the IaC-Eval benchmark, whose templates are written in Terraform. The evaluation presents inherent methodological constraints due to fundamental differences between CloudFormation and Terraform architectures as discussed in Section II-B. Consequently, we limit our evaluation to syntax validation, aligning with IaC-Eval’s original assessment methodology while acknowledging this necessary constraint. We select Claude-3.5 based on its superior performance in our primary experiments.
We observe that IaCGen successfully generalizes to Terraform syntax generation. IaCGen achieves 79.7% passItr@1 accuracy, and 100% passItr@7 accuracy, and uses an average of 1.58 iterations on IaC-Eval. In contrast, IaCGen achieves 100% passItr@5 accuracy and uses an average of 2.53 iterations to generate syntactically valid CloudFormation templates on DPIaC-Eval. This discrepancy likely stems from DPIaC-Eval containing more complex problems compared to IaC-Eval, as evidenced by our benchmark’s higher average lines of code and higher number of average iterations required. Overall, IaCGen successfully guides LLM in generating syntactically correct templates in both IaC languages, confirming its generalizability across mainstream IaC languages.
VII-B Lesson Learnt
@@
SE-Inspired LLM-based Framework. Our findings suggest that iterative, feedback-driven approaches grounded in principles of Software Engineering (SE) yield superior outcomes in IaC template generation. Rather than employing heavy prompting engineering or relying on task-specific heuristics, our framework systematically incorporates iterative feedback to guide the model across multiple refinement stages. This iterative structure mirrors conventional SE practices, such as progressive refinement, requirement validation, and configuration testing. The alignment with these established methodologies not only improves model outputs but also enhances interpretability and developer trust. Heuristics from SE may serve as a robust and generalizable research paradigm for broader AI4SE tasks, bridging the gap between LLM capabilities and practical software engineering workflows.
Vulnerability in IaC Templates. As demonstrated in our evaluation, current models, even when guided by structured frameworks like IaCGen, exhibit limited sensitivity to security concerns, potentially leading to vulnerabilities such as data breaches and service outages. For researchers, this underscores the need for advancing methods that prioritize secure IaC generation. For practitioners, it is critical to conduct security validation and expert oversight should remain integral to the deployment pipeline.
VII-C Limitation
While we believe this study makes substantial contributions into the IaC template generation, several limitations should be acknowledged. Our evaluation includes models available at the time of writing, but the rapid pace of LLM development means that newer models may show different performance characteristics. Although our benchmark focuses on CloudFormation and we evaluate the generalizabiltiy, with the similar methodology, it can easily extend to other IaC languages such as Azure Resource Manager for Azure, Google Infrastructure Manager for Google and Terraform for multi-cloud environments. As this is the first benchmark and dataset focused on IaC deployment evaluation, we chose CloudFormation because its nature supports in deployability test. Our benchmark covers 153 tasks across 58 most commonly used AWS services, but it may not capture the full diversity of real-world infrastructure requirements. Additionally, our difficulty categorization, while systematic, may not perfectly align with all organizations’ perspectives on complexity.
VIII Related Work
VIII-A LLMs for Code Generation.
LLMs have shown impressive capabilities in general-purpose code generation across various benchmarks [60, 61]. For example, Chen et al. [13] introduce HumanEval for evaluating Python code generation, while Yu et al. [62] developed Spider for assessing SQL generation performance. More recently, Liao et al. [15] proposed a context-aware LLM framework to enhance code synthesis. However, existing work primarily targets conventional programming tasks and lacks focus on infrastructure-as-code (IaC), which involves higher structural complexity and operational constraints. In contrast, our work focuses on IaC generation and introduces a domain-specific benchmark that emphasizes both syntactic quality and deployability.
VIII-B Feedback Mechanism of LLMs.

### v2

@@ 6.2. RQ2: What are the primary barriers of generating deployable IaC templates?
(1) Missing Value. The most prevalent (14.70%) error across models is the deployment error: “An error occurred (ValidationError) when calling the CreateStack operation: Parameters: [xx] must have values.” It happens when LLMs define parameters without providing corresponding values, which is particularly common with the KeyName, VPC, and Subnet parameters. Sometimes, models deliberately leave parameters undefined when information is not explicitly provided in the prompt (e.g., email addresses for SNS notifications). This error also appears to be models’ assumption that resources have already existed in the cloud environment, leading them to create placeholders for users to manually fulfill. Among all models, Claude-3.7 demonstrates the lowest occurrence rate of this error and the highest success rate in resolving it through iterative feedback. This pattern could stem from the prevalence of such template structures in the GitHub repositories as we observe.
@@
(3) Null Substitution.‘Fn::Sub’ is not needed because there are no variables” is the third most common (3.09%) error. This error typically occurs in the UserData attributes of “AWS::EC2::Instance” and “AWS::EC2::LaunchTemplate”, where the models employ “!Sub $|$ despite the absence of variables that require substitution. This pattern likely results from large-scale training corpus that contain numerous examples using this function with UserData, leading models to adopt it as a standard practice regardless of necessity.
@@
(4) Unnecessary Whitespace.Line xx: too many spaces inside brackets” accounts for 2.92% of all errors. This issue disproportionately affects o3-mini, accounting for 75% of all instances across models. The error typically manifests itself in constructs like “!Select [ 0, !GetAZs ‘’ ]” (Note the whitespaces on the left of ‘0’ and ‘]’) instead of the correct way of “!Select [0, !GetAZs ‘’]”. o3-mini consistently inserts extraneous whitespace adjacent to brackets. This error is particularly resistant to correction, often requiring explicit prompt rather than general instructions to resolve.
@@
(5) Arbitrary Default Value.Parameter validation failed: parameter value xx for parameter name xx does not exist” denotes the error that occurs when models attempt to resolve missing parameter values by providing arbitrary defaults. Claude-3.5, Claude-3.7, and DeepSeek-R1 demonstrate superior resistance in this category, averaging only 6 occurrences per model compared to approximately 27 occurrences for other models. This suggests that these models hold a more comprehensive understanding of the appropriate default values for the AWS resource parameters.
@@
Answer to RQ2: Most of errors occur during syntax stage, slightly more than those identified during deployment errors. The most common error type is Missing Value, and the top five error types account for around 30% of the total failure instances. Different models exhibit a distinct error pattern, and Claude models show the best performance in handling most of the common errors.
6.3. RQ3: How effective is human-in-the-loop feedback strategy for IaCGen?
@@
Human-in-the-loop feedback has emerged as a crucial strategy across machine learning and LLM domains (mosqueira2023human; zanzotto2019human; xiao2023llm), demonstrating significant improvements in complex scenarios especially in SE tasks. Enlightened by this, we incorporate human feedback into our framework to explore its capability on tackling unique challenges of IaC template generation. This level of feedback simulates the guidance that a senior cloud engineer might provide to a junior developer during code review, testing whether LLMs can effectively utilize and incorporate direct expert instructions.
@@
To comprehensively evaluate the effectiveness, we include human feedback in addition to the general and detailed feedback mentioned in Section 4.3. Human feedback represents the highest level of assistance in the three-tier feedback hierarchy, providing explicit correction guidance such as: “Add CidrBlock: 10.0.0.0/16 to the VPC resource” or “Replace the AMI ID with placeholder for correct AMI ID value.” A cloud engineer is asked to craft the human feedback based on the reference template and error messages. We assign the same four attempts for the human feedback as the detailed feedback with similar rationale. Therefore, the maximum attempts for each stage (Format Verification, Syntax Checking, and Live Deployment) will be ten iterations.
Figure 9. PassItr@n scores for evaluated models. Transparent lines refer to the result with general and detailed feedback. Solid lines refer to the results additionally with human feedback.
@@
Answer to RQ3: Human feedback provides consistent improvements across all evaluated models, with gains ranging from 2.6% to 37.9% percentage points. All models achieve a performance of over 90% passItr@25, and the Claude models can achieve 98% passItr@25.
6.4. RQ4: How trustworthy are the IaC templates generated by IaCGen?
User Intent Matching. User requirements for IaC templates are often ambiguously specified due to the inherent complexity and technical nature of cloud infrastructure. To evaluate how well the generated IaC templates align with the intended infrastructure requirements, we employ an intent match evaluation. Specifically, based on the DPIaC-Eval benchmark, a cloud engineering expert formulates intent specifications using Checkov’s custom policy (checkov-custom-policy), comprising two elements:
  • Required Resources: the core infrastructure components that must be presented in the template.
  • Required Attributes: the essential configuration parameters associated with these resources.
We simply use the coverage (%) to denote the level of user intent matching, i.e., how many required resources or required attributes are included in the generated templates. Table 5 shows the varying performance among models on capturing the intention of users. We observe a relatively sound capability on generating the required resources, with an average of 58.8% templates that include all specified resources across all models. In contrast, attribute-level matching presents a greater challenge, with models achieving an average of only 40.5% compliance in generating all required attributes with corresponding values. Only 25.2% of the templates satisfy both the resource-level and attribute-level requirements across all models. Claude-3.7 demonstrates superior overall performance, yet only 31.4% of its generated IaC templates fully satisfy the user intent.
Security. Security of IaC templates represents a critical concern in cloud computing, as insecure configurations can lead to data breaches, unauthorized access, and compliance violations (drosos2024your). To evaluate the security posture of LLM-generated IaC templates that passed deployability evaluation, we perform the security analysis by employing Checkov (checkov), an popular open source policy-as-code tool for IaC compliance scans (minna2025analyzing) with common industry standards such as the Center for Internet Security (cis) and AWS Foundations Benchmark (cis-aws). We define three distinct compliance metrics:
  • Policy Pass Rate (%): The percentage of security policies satisfied across all security policies checked by Checkov.
  • Unfiltered Compliance Rate (%): The proportion of generated templates that pass all applicable Checkov security policy checks, irrespective of policy applicability.
  • Filtered Compliance Rate (%): To account for Checkov’s limited policy coverage, this metric excludes templates to which no relevant security policies can be applied.
Table 4. User intent performance (%) of the generated templates at maximum iterations across models. Resource stands for resource level intent match and Attribute stands for attribute level intent match.
Model Resource Attribute Resource & Attribute
@@ 6.4. RQ4: How trustworthy are the IaC templates generated by IaCGen?
Table 5. Security performance (%) of the generated templates at maximum iterations across models. Higher numerical values indicate stronger security. Compl. stands or Compliance.
Model Policy Pass Unfiltered Compl. Filtered Compl.
@@
Table 5 shows the security compliance of generated IaC templates, and it reveals an alarmingly low security compliance rate overall. In terms of security policies, 75.3% of the examined security policies are passed on average. Surprisingly, at template-level, only 20.2% unfiltered compliance and 8.4% filtered compliance are obtained. This apparent contradiction suggests that, while models generally adhere to most security rules, critical failures in at least one policy consistently invalidate entire templates. DeepSeek-V3 demonstrates the best performance in both unfiltered and filtered compliance (24.2% and 11.5%). Claude-3.5 ranks the lowest (17.6%) in unfiltered compliance and GPT-4o shows the weakest in filtered compliance (6.1%). Further analysis reveals 55.1% of template failures originate from three AWS services: S3 Bucket, Lambda Function, and EC2 SecurityGroup. These findings indicate a critical threshold effect. While models successfully implement most security requirements (>>70.0% policy pass), failures in 1-2 policies consistently invalidate entire templates. The discrepancy between high policy-level compliance and catastrophic template-level failure highlights fundamental challenges in the ability of LLMs to generate secure IaC templates.
@@
Answer to RQ4: The trustworthiness of LLM-generated IaC templates remains moderate. Only 25.2% generated templates fully satisfy the user intent, with reasonable resource identification (58.8%) but weaker attribute fulfillment (40.5%). Security analysis shows that despite a 75.3% policy pass rate, only 8.4% of the templates achieved full compliance.
7. Discussion
7.1. Security Policy Violations in IaC Templates
As demonstrated in our RQ4 evaluation, current models exhibit limited sensitivity to security concerns, potentially leading to vulnerabilities such as data breaches and service outages. Our security analysis identified 70 distinct policy violations across all generated IaC templates, with a severity distribution of 36 low-severity, 17 medium-severity, 7 high-severity, and 10 informational-level violations according to Prisma Cloud classifications (prisma_cloud). We further analyze the top 10 most frequently violated policies, revealing one medium-severity, seven low-severity, and two informational-level violations. This distribution indicates that LLMs not only fail to implement security best practices but also potentially introduce critical vulnerabilities.
The medium-severity policy: “AWS SNS topic has SSE disabled”, reflects one systematic security oversight where LLMs fail to enable server-side encryption for message queuing services (medium_severity_policy) in generated templates. This omission exposes sensitive data in transit and at rest, violates data protection principles, and potentially compromises confidential information processed through messaging systems. In addition, violation of the low-severity policy: “AWS security groups allow ingress from 0.0.0.0/0 to port 80”, can expose AWS resources to potential security threats (low_severity_policy). Notably, low-severity does not mean it necessarily causes low security impact (dahabiyeh2021ignorance; low_severity_impact). Although classified as low-severity, this configuration creates a direct attack vector by allowing traffic from all IP addresses to the web service port. Such misconfiguration can facilitate reconnaissance attacks, denial-of-service attempts, and unauthorized access to application endpoints (mathew2014study).
The security policy violations could stem from the problematic training data. Analysis of publicly available IaC templates reveals that many contain insecure default configurations, such as security groups configured with “0.0.0.0/0” IP ranges as placeholder values (public_problematic_template_security_group), or SNS topics deployed without encryption keys for operational simplicity (public_problematic_template_SNS). LLMs appear to be pre-trained and perpetuate these problematic patterns without understanding their security implications.
For researchers, the security policy violations underscore the need to develop AI systems that integrate security considerations into their core design. Our work demonstrates the necessity for security-focused evaluation metrics in AI4SE research. Future benchmarks could incorporate multi-dimensional security assessments, including not only static policy compliance but also dynamic security analysis and threat modeling capabilities. Furthermore, practitioners should implement robust validation mechanisms before adopting LLM-generated IaC templates, to prevent security vulnerabilities from reaching production environments.
7.2. Human Feedback: Automation Potential
Even though the Claude models can achieve satisfying results (94.15% passItr@15) without human intervention, results show human feedback can enhance the overall performance. To explore the possibility on automation human feedback, we conduct a post-hoc analysis. We surprisingly find that most of human feedback could be derived from publicly available sources, such as AWS documentation (aws-doc), Stack Overflow (stackoverflow) discussions, and technical blogs. Three representative cases demonstrate how human feedback can be obtained from crowd knowledge:
S3 Bucket Naming Conflicts. GPT models commonly assign generic bucket names, causing deployment failures due to S3’s global uniqueness requirement. The original human feedback is: “S3 bucket names must be globally unique across all AWS accounts. Consider adding a unique identifier such as: random string, organization prefix, timestamp, or environment identifier to ensure uniqueness”, which guides the model to generate unique names following established naming conventions. This principle directly corresponds to the best practices documented in technical resources (bucket_naming_solution; bucket_naming_guide), which explicitly stipulate name uniqueness (“Bucket names must be unique across all AWS accounts in all of the AWS”) and solutions (“Include a Randomized Suffix or Prefix”).
Outdated AMI References. LLMs frequently provide invalid AMI identifiers for EC2 instances probably caused by biased and outdated training data, as AMI changes and is different between regions. Human feedback directs models to refer to AWS Systems Manager Parameter Store for dynamic AMI resolution. This approach follows established instructions documented in official AWS blogs (ami_solution), which demonstrate programmatic AMI lookup using the latest Amazon Linux AMI parameters (“Querying for the latest AMI using public parameters”).
Service-Specific Configuration Constraints. Complex AWS service interdependencies often result in attribute-level errors, such as the constraint “DbSubnetGroupName should not be specified for read replicas created in the same region as the master.” Human feedback resolves this by explicitly instructing LLMs to remove the conflicting attribute. Similarly, the solution is directly related to the AWS documentation (dbinstance_solution), which explicitly describes this constraint in the DbSubnetGroupName specification (“The specified DB subnet group must be in the same AWS Region…”).
Our empirical observation suggests that future iterations of IaCGen could achieve human-in-the-loop performance through LLM-based knowledge retrieval mechanisms (e.g., agentic framework, Retrieval-Augmented Generation) while maintaining full automation.
7.3. DevOps-Inspired LLM-based Framework
@@
Our findings suggest that iterative feedback-driven approaches grounded in common software development principles such as CI/CD, Agile, and Scrum yield superior outcomes in IaC template generation. Rather than employing heavy prompting engineering or relying on task-specific heuristics, our framework systematically incorporates iterative feedback to guide the model across multiple refinement stages. This iterative structure mirrors conventional SE practices, such as progressive refinement, requirement validation, and configuration testing.
The alignment with these established methodologies not only improves model outputs, but also enhances interpretability and developer trust. These DevOps-derived heuristics may serve as a robust and generalizable research paradigm for broader AI4SE tasks, bridging the gap between LLM capabilities and practical software engineering workflows. The success of this approach suggests that future AI4SE frameworks could focus on iterative refinement over single-shot generation for practical performance enhancement, particularly for complex, multi-constraint problems.
7.4. Threats to Validity
While we believe this study makes substantial contributions to LLM for IaC generation, several limitations should be acknowledged. First, our evaluation includes models available at the time of writing. The rapid evolution of LLMs may yield different performance patterns with newer releases. However, our framework’s model-agnostic design ensures compatibility with future developments. To facilitate reproducibility, we document the specific model and their cut-off date in our replication package. Second, our benchmark covers 153 scenarios across 58 commonly used AWS services, representing substantial coverage of typical infrastructure patterns. However, highly specialized configurations may not be fully captured. Also, our difficulty categorization, while systematic, may not align perfectly with all organizational perspectives on complexity. Lastly, while a comprehensive evaluation across all IaC languages would be ideal, the absence of standardized benchmarks and the substantial effort required to create such domain-specific evaluation datasets limits the broader assessment. However, our demonstrated generalizability to Terraform and IaCGen’s modular architecture enable straightforward extension to other IaC ecosystems.
8. Related Work
8.1. Feedback Mechanism of LLMs